# Tackling the Accuracy-Interpretability Trade-off in a Hierarchy of Machine Learning Models for the Prediction of Extreme Heatwaves

### v1

@@ Abstract
When performing predictions that use Machine Learning (ML), we are mainly interested in performance and interpretability. This generates a natural trade-off, where complex models generally have higher skills but are harder to explain and thus trust. Interpretability is particularly important in the climate community, where we aim at gaining a physical understanding of the underlying phenomena. Even more so when the prediction concerns extreme weather events with high impact on society.
@@
More precisely, we use models that range from a global Gaussian Approximation (GA) to deep Convolutional Neural Networks (CNNs), with the intermediate steps of a simple Intrinsically Interpretable Neural Network (IINN) and a model using the Scattering Transform (ScatNet). Our findings reveal that CNNs provide higher accuracy, but their black-box nature severely limits interpretability, even when using state-of-the-art Explainable Artificial Intelligence (XAI) tools. In contrast, ScatNet achieves similar performance to CNNs while providing greater transparency, identifying key scales and patterns in the data that drive predictions.
This study underscores the potential of interpretability in ML models for climate science, demonstrating that simpler models can rival the performance of their more complex counterparts, all the while being much easier to understand. This gained interpretability is crucial for building trust in model predictions and uncovering new scientific insights, ultimately advancing our understanding and management of extreme weather events.
The purpose of this work is to test increasingly complex machine learning models on the task of forecasting extreme heatwaves and explain their prediction. This is important to quantify what additional information the more complex models are able to capture. We find that answering this question with a black-box model and explainability techniques is not efficient. Indeed, the explanations are mainly qualitative and don’t add much to what was known from the simplest linear models. On the other hand, by using an inherently interpretable architecture design, we are able to precisely quantify the sources of additional information while maintaining the same predictive skill of the black-box model.
@@ 1 Introduction
Extreme events are responsible for the major part of weather and climate related damages; they have already become one of the most visible hallmarks of climate change over the past decades, and projections indicate that their detrimental impact will keep increasing in the future (Seneviratne et al., 2012). Heatwaves, in particular, caused significant increase in mortality in 2003 over Western Europe (Fouillet et al., 2006), in 2010 in Russia (Barriopedro et al., 2011) and in 2021 in Canada (Henderson et al., 2022). Moreover, the impact of extreme heatwaves extends to losses in the agricultural sector, stress on the power grid, increased wildfire hazards, as well as the endangerment of ecosystems (Seneviratne et al., 2021). It is thus of paramount importance to improve our understanding of these events and to be able to accurately forecast them as early as possible.
@@
Another approach to climate prediction is that of statistical models. Although such models have existed for a long time, they have recently undergone quick developments due to the recent explosion of computing power and amount of available data which has led to the widespread applications of Machine Learning to the whole field of weather and climate sciences (Bracco et al., 2024). In particular, the prediction problem has been addressed with tools ranging from simple neural networks dedicated to a specific prediction task such as heatwave (Asadollah et al., 2021; Khan et al., 2021) or El Niño index (Petersik and Dijkstra, 2020) prediction, to full foundation models that forecast the global weather with accuracy comparable to state of the art numerical simulations (Bi et al., 2023; Lam et al., 2023; Nguyen et al., 2023). Focusing on extreme events, different types of methods have been applied for heatwave prediction over different regions. For example, McKinnon et al. (2016); Vijverberg et al. (2020); Miller et al. (2021) have used regression techniques to investigate extreme heat precursors over the US, outperforming dynamical models on timescales up to four weeks. Similarly, Weirich-Benet et al. (2023) have used linear regression and random forests to forecast heatwaves over Central Europe multiple weeks in advance. Convolutional neural networks (CNNs) have been applied by Chattopadhyay et al. (2020) and by Jacques-Dumas et al. (2021); Miloshevich et al. (2023a) to the problem of heatwave classification over the US and France, respectively, again exhibiting skill deep into the S2S range.
In spite of these successes, statistical prediction for rare events still suffers from two major shortcomings. First, these models require vast amounts of data and extensive resources to be trained, especially the more complex architectures such as deep neural networks which have been increasingly used. The problem is particularly acute for extreme events, as the events with the highest impact are the rarest, which means there are very few of them in observational records and properly simulating them with climate models is expensive (Miloshevich et al., 2023a; Ragone et al., 2018). Moreover, these models essentially behave as black boxes; hence, regardless of their performance, practitioners may struggle to trust them, as they might be making the right prediction for the wrong reason, creating a natural trade-off between performance and interpretability (Rudin, 2019). In addition to consistency with scientific understanding of the underlying phenomena, gaining some degree of interpretability about the predictions would come with other benefits, such as the possibility to improve the models further, to help disentangling correlation and causality, and even to identify unknown mechanisms.
@@
However, they are fundamentally flawed as the explanations provided only approximate the true prediction of the model (Montavon et al., 2018; Yang et al., 2024) and are either understandable but simplistic or accurate but hard to analyze (Murdoch et al., 2019).
Even worse, sometimes different explanation methods can yield contrasting explanations (Mamalakis et al., 2022a, b).
@@
In this work, we address the trade-off between prediction accuracy and interpretability in the case of heatwave prediction over France. Recent work showed that regularized linear regression can be very effective while remaining interpretable (Mascolo et al., 2024). Here, we expand on this finding by training a hierarchy of increasingly complex machine learning models, from simple linear regression to fully fledged deep Convolutional Neural Networks. We explore two ways to build intermediate rungs in the hierarchy and the consequences in terms of the accuracy-interpretability trade-off: linear feature extraction followed by nonlinear prediction, or nonlinear feature extraction followed by linear regression (Sec. 3). We compare the performance of the four models, both with a dataset characteristic of the largest ones available from climate models, and with a much shorter one, more typical of observations (Sec. 4). Finally, we compare built-in interpretability of the two methods based on linear feature extraction with post-hoc explainability of the more complex networks (Sec. 5). A first aim is purely methodological. Additionally, we try to understand what extra information these latter models are able to capture with respect to their simpler counterparts. We show that linear regression already performs well and identifies commonly accepted heatwave drivers like persistent anticyclones and dry soil. CNNs have a higher performance, but post-hoc explanations provide little physical insight beyond linear regression. On the contrary, an interpretable variant of Scattering networks (Bruna and Mallat, 2013) achieves the same skill as the CNNs, but pinpoints much more precisely the source of improvement with respect to the linear regression baseline, finding extra predictability in sub-synoptic oscillations in the geopotential height field.
@@ 2.2 Heatwave amplitude
If is the temperature anomaly, we define the heatwave amplitude at as
where is the heatwave duration in days and is the geographical region of interest. In this work, will be the region of France, which has a size of roughly and thus sits nicely at the scale of cyclones and anticyclones, which is the relevant one for large-scale atmospheric dynamics. Moreover, studying heatwaves involving a whole country can be very relevant for policymakers (Barriopedro et al., 2011; Fouillet et al., 2006). We focus here on two-week heatwaves (), as longer lasting heatwaves have a much higher impact than shorter ones (Seneviratne et al., 2012, 2021; Meehl and Tebaldi, 2004; Amengual et al., 2014).
@@ 2.3 Predictors
Indeed, while the geopotential height anomaly at the middle of the troposphere gives good insight into the wind flow (thanks to the geostrophic approximation), soil moisture acts as an important modulator of the likelihood of heatwaves at mid-latitudes, as it controls the evaporative cooling potential of the surface (Perkins, 2015; Miloshevich et al., 2023a; Benson and Dirmeyer, 2021; D’Andrea et al., 2006; Fischer et al., 2007; Hirschi et al., 2011; Lorenz et al., 2010; Rowntree and Bolton, 1983; Schubert et al., 2014; Shukla and Mintz, 1982; Stefanon et al., 2012; Vargas Zeppetello and Battisti, 2020; Zeppetello et al., 2022; Zhou et al., 2019; Vautard et al., 2007). To facilitate the work of neural networks, each grid point of each of the two weather fields is independently standardized to zero mean and unitary standard deviation (see Section S.5 of the Supplementary Material).
Since this study is mainly methodological, we focus on , namely on predicting the average temperature of the next two weeks.
@@ 2.4 Probabilistic regression
We aim at predicting the distribution of the heatwave amplitude , given the current state of the predictors . From this probabilistic regression, we can build a classifier which computes the conditional probability of the heatwave amplitude exceeding a given threshold : . This quantity, called committor function (Lucente et al., 2022a), was used previously as the target for heat wave prediction with machine learning methods (Miloshevich et al., 2023a; Mascolo et al., 2024).
@@
where is the threshold that defines the 5% most extreme heatwaves.
@@ 3 The model hierarchy
To perform probabilistic regression, we will approximate the conditional distribution of the heatwave amplitude as a Gaussian distribution with mean and variance , as was done for instance for El Niño (Petersik and Dijkstra, 2020). We will use increasingly complex models to parameterize and . A summary of this model hierarchy is presented in Table 1.
@@ 3.1 Gaussian approximation
The simplest option after the climatology is to perform a linear regression of against . The underlying assumption of this method is that the joint distribution of and is a multivariate Gaussian (Mascolo et al., 2024), and results in a prediction with constant variance and . has the same dimensions as and is an optimal projection pattern, that condenses all the important information for heatwave prediction into the scalar index . Training is performed in one step, with
where and denote respectively expectation and variance of , computed on the training set. For high dimensional climate data, we need to regularize the projection pattern , and this is achieved by penalizing the norm of its spatial gradient , which forces to be spatially smooth. See Mascolo et al. (2024) for more details.
@@ 3.3 Scattering Network
Wavelet Transform and non-linearity: the input signal is convolved with a set of wavelet filters to extract frequency and phase information at different scales. The set of wavelet filters is arranged to tile the Fourier space, which is discretized by scales and orientations. Consequently, each wavelet filter is located in a distinct position in Fourier space, defined by a scale and an orientation . The result of the wavelet transform is then passed through a point-wise non-linear operator, namely the modulus operator.
Pooling and Aggregation: To achieve translation invariance, a downsampling operation is applied to the transformed coefficients. This helps in reducing the spatial resolution of the representation while retaining the essential information. We apply a local averaging operator (typically a Gaussian smoothing function), followed by an appropriate downsampling by a factor .
@@ 3.4 Convolutional Neural Network
The most complex model in this study is a Convolutional Neural Network (CNN). In this case, we feed the input to the neural network as a two-channel image, with the two ‘colors’ corresponding to the geopotential height and soil moisture fields, the latter being set to 0 outside France. This ‘image’ is processed by several convolutional layers, then by several fully connected ones, with the final output being and . The network is trained by gradient descent of the CRPS loss, and the hyperparameters of this model are its architecture plus learning rate and batch size.
@@ 3.5 Hyperparameter optimization
Hyperparameters of IINN, ScatNet and CNN are optimized using a Bayesian search algorithm provided by the optuna (Akiba et al., 2019) python package. The best combination is the one that gives the highest validation Binary Cross Entropy Skill (BCES) (see Equation (5)). We also tried optimizing with respect to the other metrics and the results do not change significantly.
The regularization coefficient of GA and IINN is treated separately, as it controls how smooth, and thus interpretable, the projection pattern is. At the best configuration of hyperparameters found by optuna, we systematically try logarithmically spaced values of . The skills of both GA and IINN display a broad plateau for intermediate values of , with poor performance for either too high or too low values (see Fig. S1 and S2 of the Supplementary Material). We then choose as the one that leads to the smoothest projection pattern without loss of performance.
@@ 4.1 Training on the full dataset
In Table 2, we show the performance of the hierarchy of models on the test set at the optimal values for hyperparameters. Since at the end of the k-fold cross validation process we have 5 trained models, we evaluate all of them on the test set, which allows us to obtain error bars on the skill of the networks. Note that since the training sets of the 5 folds share a significant portion of the data, while the validation sets are independent, the error mainly quantifies variance due to different initialization of the networks and validation on different datasets.
@@ 4.2 Training on a smaller dataset
When we test the trained models on the same 200-year-long test set as before, we obtain the results presented in Table 3, which show a reversal of the ranks, with the Gaussian approximation clearly outmatching all other methods. In Mascolo et al. (2024) the authors already point out the remarkable robustness of the Gaussian approximation to lack of data, but here we see that even the relatively similar method of the IINN suffers a lot from the dataset size, especially for the classification task, with a BCES comparable to the CNN’s.
The main reason for this seems to be the optimization method, as GA is trained in one step (see Equation (6)), while all other methods use stochastic gradient descent, which in a lack of data regime may be prone to overfitting, both during training and during the hyperparameter optimization phase (see Fig. S2 of the Supplementary Material). To test this hypothesis, we created a variant of ScatNet where is estimated with a direct linear regression from the scattering coefficients (minimizing mean square error in one step) and is kept constant at the value estimated by GA. And indeed, with this setup we were able to achieve skills that are a few percents better than GA (not shown here).

### v2

@@ Abstract
When performing predictions that use Machine Learning (ML), we are mainly interested in performance and interpretability. This generates a natural trade-off, where complex models generally have higher skills but are harder to explain and thus trust. Interpretability is particularly important in the climate community, especially when dealing with extreme weather events, where gaining a physical understanding of the underlying phenomena is crucial to contain impacts.
@@
We use models that range from a global Gaussian Approximation (GA) to deep Convolutional Neural Networks (CNNs), with the intermediate steps of a simple Intrinsically Interpretable Neural Network (IINN) and a model using the Scattering Transform (ScatNet). Our findings reveal that CNNs provide higher accuracy, but their black-box nature severely limits interpretability, even when using state-of-the-art Explainable Artificial Intelligence (XAI) tools. In contrast, ScatNet achieves similar performance to CNNs while providing greater transparency.
Our interpretable models highlight known drivers of extreme European heatwaves, like persistent anticyclonic anomalies and dry soil, as well as new ones, in the form of sub-synoptic geopotential height oscillations.
This study underscores the potential of interpretability in ML models for climate science, demonstrating that simpler models can rival the performance of their more complex counterparts, all the while being much easier to understand. This gained interpretability is crucial for building trust in model predictions and uncovering new scientific insights.
The purpose of this work is to test increasingly complex machine learning models on the task of forecasting extreme heatwaves and explain their prediction. This is important to quantify what additional information the more complex models are able to capture. We find that answering this question with a black-box model and explainability techniques is not sufficient. Indeed, the explanations are mainly qualitative and don’t add much to what was known from the simplest linear models. On the other hand, by using an inherently interpretable architecture design, we are able to identify sub-synoptic oscillations in the geopotential height field as the source of additional information, while maintaining the same predictive skill of the black-box model.
@@ 1 Introduction
Extreme events are responsible for the major part of weather and climate related damages; they have already become one of the most visible hallmarks of climate change over the past decades, and projections indicate that their detrimental impact will keep increasing in the future (Seneviratne et al., 2012). Heatwaves, in particular, caused a significant increase in mortality in 2003 in Western Europe (Fouillet et al., 2006), in 2010 in Russia (Barriopedro et al., 2011) and in 2021 in Canada (Henderson et al., 2022). Moreover, the impact of extreme heatwaves extends to losses in the agricultural sector, stress on the power grid, increased wildfire hazards, as well as the endangerment of ecosystems (Seneviratne et al., 2021). It is thus of paramount importance to improve our understanding of these events and to be able to accurately forecast them as early as possible.
@@
Another approach to climate prediction is that of statistical models. Although such models have existed for a long time, they have recently undergone quick developments due to the explosion of computing power and amount of available data which has led to the widespread applications of Machine Learning to the whole field of weather and climate sciences (Bracco et al., 2024). In particular, the prediction problem has been addressed with varying degrees of complexity, both in the tools used and in the tasks addressed. For instance, simple neural networks have been used to forecast one dimensional quantities such as heatwave (Asadollah et al., 2021; Khan et al., 2021) or El Niño (Petersik and Dijkstra, 2020) indices. On the other end of the spectrum, foundation models can forecast the global weather with accuracy comparable to state of the art numerical simulations (Bi et al., 2023; Lam et al., 2023; Nguyen et al., 2023). Focusing on extreme events, different types of methods have been applied for heatwave prediction over different regions. For example, McKinnon et al. (2016); Vijverberg et al. (2020); Miller et al. (2021) have used regression techniques to investigate extreme heat precursors over the US, outperforming dynamical models on timescales up to four weeks. Similarly, Weirich-Benet et al. (2023) have used linear regression and random forests to forecast heatwaves over Central Europe multiple weeks in advance. Convolutional neural networks (CNNs) have been applied by Chattopadhyay et al. (2020) and by Jacques-Dumas et al. (2021); Miloshevich et al. (2023a) to the problem of heatwave classification over the US and France, respectively, again exhibiting skill deep into the S2S range.
In spite of these successes, statistical prediction for rare events present a number of challenges (Materia et al., 2024). One of the two major shortcomings is the fact that these models require vast amounts of data and extensive resources to be trained, especially the more complex architectures such as deep neural networks which have been increasingly used. The problem is particularly acute for extreme events, as the events with the highest impact are the rarest, which means there are very few of them in observational records and properly simulating them with climate models is expensive (Miloshevich et al., 2023a; Ragone et al., 2018). The second issue is that neural networks essentially behave as black boxes; hence, regardless of their performance, practitioners may struggle to trust them, as they might be making the right prediction for the wrong reason, creating a natural trade-off between performance and interpretability (Rudin, 2019). In addition to consistency with scientific understanding of the underlying phenomena, gaining some degree of interpretability about the predictions would come with other benefits, such as the possibility to improve the models further, to help disentangle correlation and causality, and even to identify unknown mechanisms.
@@
However, they are only an approximation of the model’s true prediction (Montavon et al., 2018; Yang et al., 2024). Perhaps more importantly, different explanation methods can sometimes yield contrasting results (Mamalakis et al., 2022a, b).
@@
In this work, we address the trade-off between prediction accuracy and interpretability in the case of heatwave prediction over France. Recent work showed that regularized linear regression can be very effective while remaining interpretable (Mascolo et al., 2024; DelSole and Banerjee, 2017; Stevens et al., 2021). Here, we expand on this finding by training a hierarchy of increasingly complex machine learning models, from simple linear regression to fully fledged deep Convolutional Neural Networks. We explore two ways to build intermediate rungs in the hierarchy and the consequences in terms of the accuracy-interpretability trade-off: linear feature extraction followed by nonlinear prediction, or nonlinear feature extraction followed by linear regression (section 3). We compare the performance of the four models, both with a dataset characteristic of the largest ones available from climate models, and with a much shorter one, more typical of observations (section 4). Finally, we compare built-in interpretability of the two methods based on linear feature extraction with post-hoc explainability of the more complex networks (section 5). A first aim is purely methodological. Additionally, we try to understand what extra information these latter models are able to capture with respect to their simpler counterparts. We show that linear regression already performs well and identifies commonly accepted heatwave drivers like persistent anticyclones and dry soil. CNNs have a higher performance, but post-hoc explanations provide little physical insight beyond linear regression. On the contrary, an interpretable variant of Scattering networks (Bruna and Mallat, 2013) achieves the same skill as the CNNs, but pinpoints much more precisely the source of improvement with respect to the linear regression baseline, finding extra predictability in sub-synoptic oscillations in the geopotential height field.
@@ 2.2 Heatwave amplitude
We define the heatwave amplitude at as
where is the heatwave duration in days, is the geographical region of interest, and is the temperature anomaly, defined by subtracting the multi-year mean of temperature of each day.
In this work, will be the region of France, which has a size of roughly and thus sits nicely at the scale of cyclones and anticyclones, which is the relevant one for large-scale atmospheric dynamics. Moreover, studying heatwaves involving a whole country can be very relevant for policymakers (Barriopedro et al., 2011; Fouillet et al., 2006). We focus here on two-week heatwaves (), as longer lasting heatwaves have a much higher impact than shorter ones (Seneviratne et al., 2012, 2021; Meehl and Tebaldi, 2004; Amengual et al., 2014).
Note that since contains a 14-day running mean, the heatwave amplitude of consecutive days will be highly correlated.
@@ 2.3 Predictors
Indeed, while the geopotential height anomaly at the middle of the troposphere gives good insight into the wind flow (thanks to the geostrophic approximation), soil moisture acts as an important modulator of the likelihood of heatwaves at mid-latitudes, as it controls the evaporative cooling potential of the surface (Perkins, 2015; Miloshevich et al., 2023a; Benson and Dirmeyer, 2021; D’Andrea et al., 2006; Fischer et al., 2007; Hirschi et al., 2011; Lorenz et al., 2010; Rowntree and Bolton, 1983; Schubert et al., 2014; Shukla and Mintz, 1982; Stefanon et al., 2012; Vargas Zeppetello and Battisti, 2020; Zeppetello et al., 2022; Zhou et al., 2019; Vautard et al., 2007). To facilitate the work of neural networks, each grid point of each of the two weather fields is independently standardized to zero mean and unitary standard deviation (see section S.5 of the Supplementary Materials).
Since this study is mainly methodological, we will focus on a single value of the lead time . For simplicity, we chose , which means we’ll be predicting the average temperature of the coming two weeks.
@@ 2.4 Probabilistic regression
We aim at predicting the distribution of the heatwave amplitude , given the current state of the predictors . From this probabilistic regression, we can build a classifier which computes the conditional probability of the heatwave amplitude exceeding a given threshold : . This quantity, called the committor function (Lucente et al., 2022a), was used previously as the target for heat wave prediction with machine learning methods (Miloshevich et al., 2023a; Mascolo et al., 2024).
@@
where is the threshold that defines the 5% most extreme heatwaves over the training set. The index iterates through all daily snapshots of data either in the 200-year-long test set, the five 640-year-long training sets or the five 160-year-long validation sets.
@@ 3 The model hierarchy
To perform probabilistic regression, we will approximate the conditional distribution of the heatwave amplitude as a Gaussian distribution with mean and variance , as was done for instance for El Niño (Petersik and Dijkstra, 2020). We will use increasingly complex models to parameterize and . A summary of this model hierarchy is presented in Tab. 1.
@@ 3.1 Gaussian approximation
The simplest option after the climatology is to perform a linear regression of against . The underlying assumption of this method is that the joint distribution of and is a multivariate Gaussian (Mascolo et al., 2024), and results in a prediction with linear mean and constant standard deviation . has the same dimensions as and is an optimal projection pattern, that condenses all the important information for heatwave prediction into the scalar index . Training is performed in one step, with
where and denote respectively expectation and variance of , computed on the training set. For high dimensional climate data, we need to regularize the projection pattern , and this is achieved by penalizing the norm of its spatial gradient , which forces to be spatially smooth. is the residual variance of once we remove the linear trend . See Mascolo et al. (2024) for more details.
@@ 3.3 Scattering Network
Wavelet Transform and non-linearity: the input signal is convolved with a set of wavelet filters to extract frequency and phase information at different scales. The set of wavelet filters is arranged to tile the Fourier space, which is discretized by scales and orientations. Consequently, each wavelet filter is located in a distinct position in Fourier space, defined by a scale and an orientation . The result of the wavelet transform is then passed through a point-wise non-linear modulus operator.
Pooling and Aggregation: To achieve local translation invariance, a downsampling operation is applied to the transformed coefficients. This helps in reducing the spatial resolution of the representation while retaining the essential information. We apply a local averaging operator (typically a Gaussian smoothing function), followed by an appropriate downsampling by a factor .
@@ 3.4 Convolutional Neural Network
The most complex model in this study is a Convolutional Neural Network (CNN). In this case, we follow a similar setup to Miloshevich et al. (2023a), where we feed the input to the neural network as a two-channel image, with the two ‘colors’ corresponding to the geopotential height and soil moisture fields, the latter being set to 0 outside France. This ‘image’ is processed by several convolutional layers, then by several fully connected ones, with the final output being and . The network is trained by gradient descent of the CRPS loss, and the hyperparameters of this model are its architecture plus learning rate and batch size.
@@ 3.5 Hyperparameter optimization
Hyperparameters of IINN, ScatNet and CNN are optimized using a Bayesian search algorithm provided by the optuna (Akiba et al., 2019) python package. The best combination is the one that gives the highest validation Binary Cross Entropy Skill (BCES) (see Eq. (5)). We also tried optimizing with respect to the other metrics and the results do not change significantly.
The regularization coefficient of GA and IINN is treated separately, as it controls how smooth, and thus interpretable, the projection pattern is. At the best configuration of hyperparameters found by optuna, we systematically try logarithmically spaced values of . The skills of both GA and IINN display a broad plateau for intermediate values of , with poor performance for either too high or too low values (see Figs. S1 and S2 of the Supplementary Materials). We then choose as the one that leads to the smoothest projection pattern without loss of performance.
@@ 4.1 Training on the full dataset
In Tab. 2, we show the performance of the hierarchy of models on the test set at the optimal values for hyperparameters. Since at the end of the k-fold cross validation process we have 5 trained models, we evaluate all of them on the test set, which allows us to obtain error bars on the skill of the networks. Note that since the training sets of the 5 folds share a significant portion of the data, while the validation sets are independent, the error mainly quantifies variance due to different initialization of the networks and validation on different datasets.
@@ 4.2 Training on a smaller dataset
When we test the trained models on the same 200-year-long test set as before, we obtain the results presented in Tab. 3, which show a reversal of the ranks, with the Gaussian approximation clearly outmatching all other methods. In Mascolo et al. (2024) the authors already point out the remarkable robustness of the Gaussian approximation to lack of data, but here we see that even the relatively similar method of the IINN suffers a lot from the dataset size, especially for the classification task, with a BCES comparable to the CNN’s.
The main reason for this seems to be the optimization method, as GA is trained in one step (see Eq. (6)), while all other methods use stochastic gradient descent, which in a lack of data regime may be prone to overfitting, both during training and during the hyperparameter optimization phase (see Fig. S2 of the Supplementary Materials). To test this hypothesis, we created a variant of ScatNet where is estimated with a direct linear regression from the scattering coefficients (minimizing mean square error in one step) and is kept constant at the value estimated by GA. And indeed, with this setup we were able to achieve skills that are a few percentage points better than GA (not shown here).